OpenStudy (cenaida):

Graph the line with slope 2/5 and y intercept -2

OpenStudy (ybarrap):

y=mx+b, is general equation of a line. "m" is the slope, "b" is the y-intercept. You have m=2/5 and b=-2. Any questions?

OpenStudy (ybarrap):

If you want to graph, make these substitutions then plug in any x and see what y you get, do that for 2 or 3 points, connect the dots together, and that will be your graph.

OpenStudy (cenaida):

@JuanitaM @aysian97 please help

OpenStudy (anonymous):

Okay, so basically you have a line graph. Now, the slope (or "m") is 2/5. So that's means "rise over run". It goes up 2 and over 5 spaces. If you know the y intercept is -2, then go up 2 and over 5. What is your new coordinate?

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

I'm assumed the x value was 0 when i explained that to you...just saying

OpenStudy (anonymous):

did you HAVE an x value in the question?

OpenStudy (anonymous):

@ybarrap he was right when explaining it, too

OpenStudy (cenaida):

Can u draw the graph for me

OpenStudy (anonymous):

i can try hold on...

OpenStudy (cenaida):

Ok thanks

OpenStudy (anonymous):

Okay..so the drawing is really not working for me! I tried like 3 times and i can't get it to look right. I'm sorry

OpenStudy (cenaida):

Ok ... thanks anyways...

OpenStudy (anonymous):

you're very welcome